# An Analysis of Hyper-Parameter Optimization Methods for Retrieval Augmented Generation

### v1

@@ Abstract
Finding the optimal Retrieval-Augmented Generation (RAG) configuration for a given use case can be complex and expensive. Motivated by this challenge, frameworks for RAG hyper-parameter optimization (HPO) have recently emerged, yet their effectiveness has not been rigorously benchmarked. To address this gap, we present a comprehensive study involving 5 HPO algorithms over 5 datasets from diverse domains, including a new one collected for this work on real-world product documentation. Our study explores the largest HPO search space considered to date, with two optimized evaluation metrics. Analysis of the results shows that RAG HPO can be done efficiently, either greedily or with iterative random search, and that it significantly boosts RAG performance for all datasets. For greedy HPO approaches, we show that optimizing models first is preferable to the prevalent practice of optimizing sequentially according to the RAG pipeline order.
An Analysis of Hyper-Parameter Optimization Methods for Retrieval Augmented Generation
@@ 1 Introduction
Despite ongoing advances in applying Large Language Models (LLMs) to real-world use cases, ensuring that LLMs have access to required knowledge to resolve user queries is a challenge. This has led to the Retrieval-Augmented Generation (RAG) paradigm, where a retrieval system provides the generative LLM with a context of grounded information (Lewis et al., 2020; Huang and Huang, 2024; Gao et al., 2024; Wang et al., 2024b).
The popularity of RAG is largely thanks to its modular design. By relying on a dedicated retrieval component, RAG solutions focus LLMs on grounded data, reducing the likelihood of dependence on irrelevant preexisting knowledge. These solutions also allow full control over which data sources to pull information from for each user question. The data sources can be refreshed regularly, thus keeping the system up-to-date.
@@
The modularity of RAG also means that practitioners are faced with a wide array of decisions when designing their RAG pipelines. One such choice is which generative LLM to use; other choices pertain to parameters of the retrieval system, such as how many items to retrieve per input question, how to rank those items, and so forth.
@@
Furthermore, evaluating even a single RAG configuration is costly in terms of time and funds: the embedding step as part of corpus indexing is compute-intensive; generating answers using LLMs is also a demanding task, especially for large benchmarks; and evaluation with LLMaaJ adds another costly round of inference.
Therefore, exploring and evaluating the full exponential search space of retrieval and generation parameters is prohibitively expensive; and at the same time, suboptimal parameter choices may significantly harm performance. Moreover, assuming that one finds a set of RAG parameters that work well on a particular dataset, there is no guarantee that the same parameters will perform well on other datasets or domains, or even on the same dataset as the source documents or user queries undergo data drift over time, and as the available technologies (e.g., generative and embedding models) evolve.
@@
One approach for addressing these issues is to use automated hyper-parameter optimization (HPO) for RAG (Fu et al., 2024a; Kim et al., 2024). The general idea of such approaches is to conduct experiments on a limited set of RAG configurations, and recommend a configuration based on the results. Experiments can be performed in various ways, ranging from established HPO algorithms to a naive random sampling of configurations to test. Importantly, although HPO frameworks are already gaining ground in RAG use-cases, their effectiveness has not been rigorously tested.
In this work we study the question of how well HPO algorithms work in practice for RAG use cases. To this end, we conduct a comprehensive study involving $5$ HPO algorithms: Tree-Structured Parzen Estimators (TPE) (Watanabe, 2023), $3$ greedy optimizations, and random search. Our experiments explore a large search space of $162$ possible RAG configurations, formed from $5$ parameters of retrieval and generation. To our knowledge, this is more than twice the size of the largest RAG search space previously explored for HPO. [^1] Fu et al. (2024b) explored a search space of $75$ possible configurations formed from $3$ parameters. See §2. Our study also takes into account both traditional lexical metrics and more recent LLMaaJ approaches.
Our experiments span a diverse set of $5$ RAG question-answering datasets from different domains: arXiv articles on ML (Eibich et al., 2024), Biomedical (Krithara et al., 2023), Wikipedia (Rosenthal et al., 2024; Smith et al., 2008) and a new product documentation dataset we collect and release as open-source.
We emphasize that in all our experiments, we select RAG configurations based on their performance on a development set, but report their evaluation on held out test data. This accurately models the real-world scenario where a dataset is provided up front for experimentation but unseen user questions arrive after deployment.
@@
The main contributions of this work are: (i) benchmark results over the largest RAG search space yet to be explored (see above), showing that RAG HPO can be done efficiently, either greedily or with simple random search, and that it significantly boosts RAG performance for all datasets; (ii) a detailed analysis of the results, exploring the connections between the optimized parameters, the dataset and the optimization objective. For greedy HPO approaches, we show that optimizing models first is preferable to the prevalent practice of optimizing sequentially according to pipeline order; (iii) a new open-source RAG dataset over enterprise product documentation. [^2] https://huggingface.co/datasets/ibm-research/watsonxDocsQA
@@ 2 Related Work
Prior work on HPO for RAG includes AutoRAG-HP Fu et al. (2024b), which explores $75$ RAG configurations induced from $3$ parameters, of which our work includes two. In comparison, our search space spans $162$ configurations, and adds the choice of a generative model, which we show to be the most dominant RAG parameter (see §6). Also, Fu et al. (2024b) report results on the same set they optimize on, hiding potential over-fitting, while our reports are on a held-out test set.
@@ 3 Experimental Setup
The architecture of different RAG pipelines is diverse, with some components common to most architectures. We focus on these common parts, which we represent by the following linear flow. Processing starts with chunking the input documents into smaller chunks, based on two predefined parameters: the size of each chunk in tokens, and the overlap between consecutive chunks. This is followed by representing each chunk with a dense vector created by an embedding model – our third parameter. The vectors are stored in a vector-database, [^6] Milvus (Wang et al., 2021) with default settings; index type is HNSW Malkov and Yashunin (2018). alongside their original text. Upon receiving a query, it is embedded too, and the Top-K ($K$ being our forth parameter) relevant chunks are retrieved (retrieval). Lastly, a prompt containing the query and the retrieved chunks is passed to a generative model (our fifth parameter) to create an answer (generation). [^7] Generation details are in Appendix F. The flow of our RAG pipeline together with the hyper parameters we used (see below) is depicted in Fig 1.
Figure 1: Our RAG pipeline with hyper-parameters.
@@ 3.1 Search space
Parameter RAG step Values
@@
Table 1: Hyper-Parameters in Explored RAG Configurations
@@ 3.2 Datasets
(ProductDocs) A new open source dataset and benchmark based on enterprise product documentation. It consists of $5534$ sub-documents generated from $1144$ pages of HTML product documentation. As with the ClapNQ dataset, the sub-documents serve as input for the pipeline. The benchmark includes $75$ QA pairs and gold document labels, of which $25$ were generated by two subject matter experts, and the rest were generated synthetically and then manually filtered and reviewed by two of the authors.
@@
Every benchmark was split into development and test sets. To keep computations tractable, the number of questions in the large benchmarks (BioASQ and ClapNQ) was limited to $1000$ for development and $150$ for test. Table 2 lists the corpora and benchmark sizes and domains. Question examples are in Appendix A.
@@
3.3 Evaluation
Retrieval Quality is evaluated using context correctness with Mean Reciprocal Rank Voorhees and Tice (2000).111111 Note that as labeling is at the document level, any chunk from the gold document is considered as a correct prediction, even if it does not include the answer to the question.
Answer correctness compares generated and gold answers, assessing overall pipeline quality, and is measured in two ways: First, a fast, lexical, implementation based on token recall (Lexical-AC), which provides a good balance between speed and quality Adlakha et al. (2024). Second, a standard LLM-as-a-Judge answer correctness (LLMaaJ-AC) implementation from the RAGAS library Es et al. (2023), with GPT4o-mini OpenAI (2024) as its backbone. This implementation performs $3$ calls to the LLM on every invocation, making it much slower and more expensive than the lexical variant.
While iterating over a benchmark, context correctness is computed per question after retrieval, and answer correctness is computed after answer generation. Averaging these scores over all benchmark instances determines the global scores of the RAG configuration on the dataset.
@@ 4 HPO Algorithms
The HPO algorithms in our experiments are iterative. At each iteration, the algorithm looks at all previously explored RAG configurations and their scores, and chooses one unexplored configuration to try next. Simulating a fixed exploration budget, after a fixed number of iterations the algorithm stops and the best performing configuration is returned. A good HPO algorithm identifies a top-performing configuration with minimal iterations.
We consider two types of HPO algorithms. The first are standard HPO algorithms not specifically tailored to a certain type of RAG pipeline. The other type of algorithms make specific assumptions on the optimized pipeline and its components. All optimize an answer correctness metric (see §3.3) unless explicitly stated otherwise below.
Our first standard HPO algorithm is the Tree-Structured Parzen Estimators (TPE) algorithm Watanabe (2023), implemented in the hyperopt library Bergstra et al. (2013). We configure it to start with $5$ random bootstrapping iterations and otherwise use the default settings. The second standard algorithm (Random) simply samples a new random RAG configuration, without replacement, at each iteration, while ignoring the scores of previously explored configurations.
The other algorithms are greedy, and require a list of search space parameters as input, ordered by their presumed impact on the RAG pipeline. They iterate the list, optimizing one parameter at a time, assuming that optimizing more impactful ones first leads to quick identification of a good configuration. When optimizing a list parameter $p$, assuming values for those before on the list were already determined, the algorithm samples random values for all following list parameters. The parameter $p$ is then evaluated with all possible values. The value yielding the best objective score is picked, and the algorithm continues to the next hyper-parameter.
The greedy algorithms differ in the order in which the hyper-parameters are optimized. One option (Greedy-M) is to order the generative and embedding models first, assuming they are more important: Generative Model, Embedding Model, Chunk Size, Chunk Overlap, Top-k. A second prevalent one (Greedy-R) is to optimize by the order of the RAG pipeline itself, first the retrieval part (still with the model first), followed similarly by the generation part: Embedding Model, Chunk Size, Chunk Overlap, Generative Model and Top-k. A third option (also prevalent) (Greedy-R-CC) uses the same order, yet optimizes the first $3$ retrieval-related parameters according to the score of a context correctness metric evaluated on the produced index, thus avoiding any LLM-based inference until all retrieval parameters are chosen. After that, the algorithm optimizes the remaining parameters with answer correctness.
@@ 5 Results
The experiments consider the search space of §3.1 and all described datasets in §3.2. The optimization objective is answer correctness, implemented with either LLMaaJ-AC or Lexical-AC (§3.3).
Every HPO algorithm ran on each development set for $10$ iterations. After each iteration, the best configuration so far was applied to the test set. This allows a per-iteration progress analysis (see below). Since all algorithms involve a random element, each run was repeated with $10$ different seeds.
Our setting was chosen to simulate a realistic use case: an HPO algorithm is executed over a benchmark (the development set) and is used to select a RAG configuration to deploy; the deployed configuration is expected to generalize to a set of unseen questions – simulated here via a test set.
To establish upper bounds on the expected performance, we also run a full grid search over the development and test sets. These upper bounds allow us to assess to what extent the different algorithms approach the optimal result. This entails evaluating all possible $162$ configurations, including $18$ different indexes, on all $5$ datasets.
@@
Figure 2 depicts the per-iteration performance of the algorithms on the test sets. Also shown (black dashed line) is the best achievable performance on test, obtained by a full grid search on that set. For all algorithms, it is enough to explore around $10$ configurations to obtain performance that is comparable to exploring all $162$ available configurations with a naive grid search (solid red lines in Figure 2). This strong result holds for both evaluation metrics.
@@
Also evident in the results is the improvement at iteration $10$ over iteration $1$. The latter, being a single random choice, may be thought of as a proxy for a naive unexperienced user choosing parameter values in a manner that is similar to random. This demonstrates the benefit of HPO for practitioners making their first steps in RAG.
Focusing on the greedy algorithms, the results imply that the order in which parameters are optimized is of great importance. Specifically, the methods that start by optimizing retrieval-related parameters (Greedy-R and Greedy-R-CC) require more iterations to identify good configurations.
The naive option of random sampling also finds a good RAG configuration after a small number of iterations. The complex TPE algorithm does not appear to add much value, choosing a configuration which is roughly equivalent in quality to random choice. This may be a result of the size of our search space, and also because 10 iterations is relatively few for TPE which in any case begins with several random iterations. Exploring larger search spaces, possibly giving up on a full grid search comparison, is a topic for further work.
@@
The above results were produced with each RAG pattern evaluated on an entire development set. While simple, this option is costly for large datasets. Research in other domains has suggested that limiting evaluation benchmarks through random sampling improves efficiency without compromising evaluation quality Perlitz et al. (2024); Polo et al. (2024). Thus, we experiment with HPO that uses a sample of the development set. To our knowledge, this is the first study of sampling in the context of RAG HPO.
@@
Following sampling, we repeat the experiments using the best HPO methods: Random, TPE, and Greedy-M. Figure 3 depicts the results when optimizing for LLMaaJ-AC, comparing all three methods when applied to the full (solid lines) or sampled (dotted) development sets (similar Lexical-AC results are in Appendix B). As can be seen, for BioASQ results remain about the same, while for ClapNQ sampling leads to identifying a suboptimal configuration after all $10$ iterations, most notably for TPE and Random.

### v2

@@ Abstract
Finding the optimal Retrieval-Augmented Generation (RAG) configuration for a given use case can be complex and expensive. Motivated by this challenge, frameworks for RAG hyper-parameter optimization (HPO) have recently emerged, yet their effectiveness has not been rigorously benchmarked. To address this gap, we present a comprehensive study involving 5 HPO algorithms over 5 datasets from diverse domains, including a new one collected for this work on real-world product documentation. Our study explores the largest HPO search space considered to date, with three evaluation metrics as optimization targets. Analysis of the results shows that RAG HPO can be done efficiently, either greedily or with random search, and that it significantly boosts RAG performance for all datasets. For greedy HPO approaches, we show that optimizing model selection first is preferable to the prevalent practice of optimizing according to RAG pipeline order.
An Analysis of Hyper-Parameter Optimization Methods
for Retrieval Augmented Generation
@@ 1 Introduction
In the Retrieval-Augmented Generation (RAG) paradigm, a generative LLM answers user questions using a retrieval system which provides relevant context from a corpus of documents (Lewis et al., 2020; Huang and Huang, 2024; Gao et al., 2024; Wang et al., 2024b). By relying on a dedicated retrieval component, RAG solutions focus LLMs on grounded data, reducing the likelihood of dependence on irrelevant preexisting knowledge.
@@
The popularity of RAG is largely thanks to its modular design, allowing full control over which data sources to pull data from and how to process that data. While advantageous, this modularity also means that practitioners are faced with a wide array of decisions when designing their RAG pipelines. One such choice is which generative LLM to use; other choices pertain to parameters of the retrieval system, such as how many items to retrieve per input question, how to rank them, and so forth.
@@
Furthermore, evaluating even a single RAG configuration is costly in terms of time and funds: the embedding step as part of corpus indexing is compute-intensive; generating answers using LLMs is also a demanding task, especially for large benchmarks; and evaluation with LLMaaJ adds another costly round of inference. Therefore, exploring and evaluating the full exponential search space of RAG parameters is prohibitively expensive; and at the same time, suboptimal parameter choices may significantly harm result quality.
@@
One approach for addressing these issues is to use automated hyper-parameter optimization (HPO) for RAG (Fu et al., 2024; Kim et al., 2024): conducting experiments on a limited set of RAG configurations, and recommending a configuration based on the results. Experiments can be performed in various ways, ranging from established HPO algorithms to a naive random sampling of configurations to test. Importantly, although HPO frameworks are gaining ground in RAG use-cases, their effectiveness has not been rigorously tested.
Figure 1: We study hyper parameter optimization over a RAG pipeline with $5$ parameters. The explored search space includes $162$ RAG configurations formed from combinations of the depicted hyper parameters.
In this work we study the question of how well HPO algorithms work in practice for standard RAG pipelines comprising retrieval and generation (see Figure 1). Our experiments explore the largest RAG search space considered to date, consisting of $162$ possible RAG configurations, derived from five parameters: chunk size and overlap, which control how documents are split, the embedding model used to encode chunks in a vector database, the number of retrieved chunks included as context when answering a question, and the generative model that produces the answer. This search space is explored with $5$ HPO algorithms: Tree-Structured Parzen Estimators (TPE) (Watanabe, 2023), $3$ greedy algorithms, and random search. Our evaluation also includes grid search over the entire search space, enabling comparison to the best possible result. The study of a larger search space induced by more complex RAG pipelines with additional parameters (e.g. agentic pipelines, adding rerankers, etc.) is computationally demanding for grid search and deferred to future work.
Our experimental setup follows the real-world scenario where a dataset is provided up front for experimentation and unseen user questions arrive after deployment. We therefore select RAG configurations based on their performance on a development set, and report their results on held out test data. To represent real use cases, we employ a diverse set of $4$ existing RAG question-answering datasets from different domains: arXiv articles on ML (Eibich et al., 2024), Biomedical (Krithara et al., 2023), Wikipedia (Rosenthal et al., 2024; Smith et al., 2008). In addition, we experiment with a new fifth dataset over the domain of product documentation, which was collected and open-sourced as part of this work. Lastly, as different applications optimize according to different dimensions (e.g. answer correctness or faithfulness), our study also takes into account different RAG objectives implemented in various ways.
@@
Our main contributions are: (i) benchmark results over the largest RAG search space to-date, showing that RAG HPO can be done efficiently, either greedily or with simple random search, and that it significantly boosts RAG performance for all datasets; (ii) a detailed analysis of the results, exploring the connections between the optimized parameters, the dataset and the optimization objective. For greedy HPO approaches, we show that the order in which the parameters are optimized is of great importance; (iii) a new open-source RAG dataset over enterprise product documentation. [^1] https://huggingface.co/datasets/ibm-research/watsonxDocsQA
@@ 2 Related Work
Prior work on HPO for RAG includes AutoRAG-HP Fu et al. (2024), which explores $75$ RAG configurations induced from $3$ parameters. In comparison, our search space spans $162$ configurations, and adds the choice of a generative model, which we show to be the most dominant RAG parameter (see §6). Importantly, Fu et al. (2024) focus on a setting with online feedback from users, and thus explore an online algorithm that iteratively updates the reward for the various RAG parameters based on small batches of queries. In contrast, our study involves offline HPO procedures which take full benchmarks as input. Also, Fu et al. (2024) report results on the same set they optimize on, hiding potential over-fitting, while our reports are on a held-out test set, simulating real-world performance on unseen user queries.
@@ 3 Experimental Setup
In our explored RAG pipeline (see Figure 1), processing starts with chunking the input documents into smaller chunks, based on two predefined parameters: the size of each chunk in tokens, and the overlap between consecutive chunks. This is followed by representing each chunk with a dense vector created by an embedding model – our third parameter. The vectors are stored in a vector-database, [^5] Milvus (Wang et al., 2021) with default settings; index type is HNSW Malkov and Yashunin (2018). alongside their original text. Upon receiving a query, it is embedded too, and the Top-K ($K$ being our fourth parameter) relevant chunks are retrieved (retrieval). Lastly, a prompt containing the query and the retrieved chunks is passed to a generative model (our fifth parameter) to create an answer (generation). [^6] Generation details are in Appendix F.
@@ 3.1 Search space
Hyper Parameter RAG step Values
@@
Table 1: The hyper parameters explored in our search space, and their values.
@@ 3.2 Datasets
(ProductDocs) A new open source dataset and benchmark based on enterprise product documentation. It consists of $5534$ sub-documents generated from $1144$ pages of HTML product documentation. As with the ClapNQ dataset, the sub-documents serve as input for the pipeline. The benchmark includes $75$ QA pairs and gold document labels, of which $25$ were generated by two subject matter experts, and the rest were generated synthetically and then manually filtered. All QA pairs were additionally reviewed by two of the authors, ensuring high data quality. Further details and examples from the datasets are in Appendix A.1.
@@
Every benchmark was split into development and test sets. To keep computations tractable, the number of questions in the large benchmarks (BioASQ and ClapNQ) was limited to $1000$ for development and $150$ for test. Table 2 lists the corpora and benchmark sizes and domains.
@@
3.3 Metrics
The following metrics were used in our experiments: Retrieval quality was measured using context correctness with Mean Reciprocal Rank Voorhees and Tice (2000).101010 Note that as labeling is at the document level, any chunk from the gold document is considered as a correct prediction, even if it does not include the answer to the question. Answer faithfulness (Lexical-FF) measures whether a generated answer remained faithful to the retrieved contexts, with lexical token precision. Answer correctness compares generated and gold answers, assessing overall pipeline quality, and is measured in two ways: First, a fast, lexical, implementation based on token recall (Lexical-AC), which provides a good balance between speed and quality Adlakha et al. (2024). Second, a standard LLM-as-a-Judge answer correctness (LLMaaJ-AC) implementation from the RAGAS library Es et al. (2023), with GPT4o-mini OpenAI (2024) as its backbone. This implementation performs $3$ calls to the LLM on every invocation, making it much slower and more expensive than the lexical variant.
Given a benchmark, all metrics were computed per-question. Averaging the per-question scores yields the overall metric score.
We note that all benchmarked HPO approaches are metric-agnostic - they are not tailored to any specific metric, nor a specific performance axis (such as answer correctness). Similarly, HPO can be applied to multiple RAG metrics at once, forming a single optimization objective.
@@ 4 HPO Algorithms
The HPO algorithms in our experiments are iterative. At each iteration, the algorithm looks at all previously explored RAG configurations and their scores, and chooses one unexplored configuration to try next. Simulating a limited exploration budget, after a fixed number of iterations the algorithm stops and the best performing configuration is returned. A good HPO algorithm identifies a top-performing configuration with minimal iterations.
We consider two types of HPO algorithms. The first are standard HPO algorithms not specifically tailored to RAG. The other type required some knowledge of the optimized RAG pipeline and its components. All optimize answer correctness (see §3.3) unless explicitly stated otherwise.
Our first standard HPO algorithm is the Tree-Structured Parzen Estimators (TPE) algorithm Watanabe (2023), using an implementation from hyperopt Bergstra et al. (2013). It was configured with $5$ random bootstrapping iterations and otherwise the default settings. The second standard algorithm (Random) samples an unexplored random RAG configuration at each iteration, while ignoring results from previous iterations.
The other type of algorithms take as input a list of search space parameters ordered by their presumed impact on the RAG pipeline. These algorithms take a greedy approach: they iterate the list, optimizing one parameter at a time, assuming that optimizing more impactful ones first leads to quick identification of a good configuration. When optimizing a parameter $p$ the algorithm assumes that values for preceding parameters were already determined. The parameter $p$ itself is evaluated with all possible values, while the values of all following parameters are chosen randomly. The value yielding the best objective score is picked for $p$, and the algorithm continues to the next hyper-parameter.
The greedy algorithms differ solely in the order in which the hyper-parameters are optimized. One option (Greedy-M) is to order the generative and embedding models first, assuming they are more important: Generative Model, Embedding Model, Chunk Size, Chunk Overlap, Top-k. A second prevalent one (Greedy-R) is to optimize by the order of the RAG pipeline itself, first the retrieval part (still with the model first), followed similarly by the generation part: Embedding Model, Chunk Size, Chunk Overlap, Generative Model and Top-k. A third option (also prevalent) (Greedy-R-CC) uses the same order, yet optimizes the first $3$ retrieval-related parameters according to the score of a context correctness metric evaluated on the produced index, thus avoiding any LLM-based inference until all retrieval parameters are chosen. After that, the algorithm optimizes the remaining parameters with answer correctness.
@@ 5 Results
The experiments consider the search space of §3.1 and all described datasets in §3.2. The optimization objective was answer correctness, implemented with either LLMaaJ-AC or Lexical-AC, or answer faithfulness with Lexical-FF (§3.3).
Our experimental setting was chosen to simulate a realistic use case: an HPO algorithm is executed over a benchmark (the development set) and the best RAG configuration is chosen for deployment; the deployed configuration is expected to generalize to a set of unseen questions – simulated here via a test set.
Specifically, every HPO algorithm ran on each development set for $10$ iterations. After each iteration, the best configuration on the development set was applied to the test set. This allows a per-iteration progress analysis (see below). Since all algorithms involve a random element, each run was repeated with $10$ different seeds.
To establish upper bounds on the expected performance, we also ran a full grid search over the development and test sets. These upper bounds allow us to assess to what extent the different algorithms approach the optimal result. This entails evaluating all possible $162$ configurations, including $18$ different indexes, on all $5$ datasets.
@@
Figure 2 depicts the per-iteration performance of the HPO methods on the test sets, when optimizing for answer correctness. Results for answer faithfulness are of the same trend, and are therefore deferred to Appendix §B. In each plot, the best achievable performance, obtained by running a full grid search on the test set and selecting the best configuration, is shown with a black dashed line.
For all algorithms, the plots demonstrate that it is enough to explore around $10$ configurations to obtain performance that is comparable to exploring all $162$ available configurations with a naive grid search (solid red lines in Figure 2). This strong result holds for both evaluation metrics.
@@
Focusing on the greedy algorithms, the results suggest that the order in which the parameters are optimized is of great importance. Specifically, methods that start by optimizing retrieval-related parameters (Greedy-R and Greedy-R-CC) require more iterations to reach good configurations.
The naive option of random sampling also finds a good RAG configuration after a small number of iterations. The complex TPE algorithm was found to be roughly equivalent in quality to random choice in our setup. This presumably stems from the size of our search space. In even larger search spaces (for example, with continuous parameters) further exploration of TPE may prove valuable.
@@
The above results were produced with each RAG pattern evaluated on the whole development set. While simple, this option is costly for large datasets. Research in other domains has suggested that limiting evaluation benchmarks by random sampling improves efficiency without compromising evaluation quality Perlitz et al. (2024); Polo et al. (2024). Thus, we experiment with HPO that uses a sample of the development set. To our knowledge, this is the first study of sampling in the context of RAG HPO.
@@
Following sampling, we repeat the experiments using the best HPO methods: Random, TPE, and Greedy-M. Figure 3 depicts the results when optimizing for LLMaaJ-AC, comparing all three methods when applied to the full (solid lines) or sampled (dotted) development sets (similar Lexical-AC results are in Appendix B). For BioASQ the results remain about the same, while for ClapNQ sampling leads to identifying a suboptimal configuration after all $10$ iterations, most notably for TPE and Random.